OpenStudy (anonymous):

Write an equation of the line that passes through points (3, 7) and (5, 11).

OpenStudy (anonymous):

The slope is 2 (for every increase in x, y increases 2). When x = 0 y = 1 (by slope, or draw it). Formula is y = 2x + 1

OpenStudy (anonymous):

thesee are the multiple choice answers. y – 7 = 5(x – 3) y – 7 = 2(x – 3) y + 7 = 5(x – 3) y + 7 = 2(x + 3)

OpenStudy (anonymous):

The answer is y - 7 = 2(x - 3)

OpenStudy (anonymous):

Could you help me with another question @anessaleyva

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

oopss

OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3

OpenStudy (anonymous):

y –8 =3(x + 8) y + 8 = 3(x + 8) y – 8 = 3(x – 8) y + 8 = 3(x – 8)

OpenStudy (anonymous):

y + 8 = 3(x - 8)

OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = –1.7x + 8.5

OpenStudy (anonymous):

17x + 10y = 85 –17x + 10y = 85 –17x + 10y = –85 17x + 10y = –85

OpenStudy (anonymous):

Sorry, I'm afraid I can't answer that one

OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = –1.7x + 8.5

OpenStudy (anonymous):

17x + 10y = 85 –17x + 10y = 85 –17x + 10y = –85 17x + 10y = –85